OpenStudy (mall2697):

Suppose you increase your walking speed from 7 m/s to 13 m/s in a period of 1 s. What is your acceleration?

OpenStudy (mangotangochick):

a = (v-u)/t a=(13-7)/1 Therefore your acceleration would be 6m/s^-2

OpenStudy (mangotangochick):

Whoops sorry I mean 6ms^-2

OpenStudy (osprey):

13m/s-6m/s=7m/s. That in 1 sec is 7m/s/s That's SOME big acceleration - gravity is 10m/s/s ! http://perendis.webs.com

OpenStudy (agent0smith):

13m/s is not exactly walking pace. That's 47 km/h. Probably faster than Usain Bolt can run at full speed.

OpenStudy (mangotangochick):

Usain Bolt top speed: (According to Google :)) 44.64kmh = 12.40ms Just about faster than Usain ;)

OpenStudy (osprey):

ouch ... getting near torn muscle territory for "non Usain"s